OpenStudy (undeaddragon26):

Find the length of c if a=62 degrees and angle b =39 degrees and b = 15.

OpenStudy (imstuck):

is 62 an angle measure?

OpenStudy (undeaddragon26):

yea

OpenStudy (imstuck):

ok I got you...let me set it up.

OpenStudy (imstuck):

First thing you need is the measure of angle C. You can find that by subtracting 39 and 62 from 180. The difference is C.

OpenStudy (undeaddragon26):

79

OpenStudy (imstuck):

Angle C is...79, right!

OpenStudy (undeaddragon26):

yea

OpenStudy (imstuck):

Now you need to tell me that you are familiar with the Law of Sines

OpenStudy (undeaddragon26):

sorta im a little rusty

OpenStudy (imstuck):

Ok, it looks like this:

OpenStudy (imstuck):

\[\frac{ sinA }{ a }=\frac{ sinB }{ b}=\frac{ sinC }{ c }\]There are 2 things you need to know about this: the lowercase letters are side lengths while the uppercase are angles, and you only need two of those identities to solve for either a missing length or missing side.

OpenStudy (imstuck):

You have angle B and the side length of b, you also have angle C now and you want side c. So that gets set up like this:

OpenStudy (imstuck):

\[\frac{ sinB }{ b}=\frac{ sinC }{ c }\]and you have three of the 4 of those needed. You have this:\[\frac{ \sin(39) }{ 15 }=\frac{ \sin(79) }{ c }\]

OpenStudy (imstuck):

To solve for side c you will cross multiply. Can you do that?

OpenStudy (undeaddragon26):

1185=39c

OpenStudy (imstuck):

here...

OpenStudy (imstuck):